# Critiques of World Models

### v2

@@ 4.3 Architecture: Autoregressive Generation is Not the Enemy
A common concern with encoder-decoder architectures (which defines an additional function $\hat{o}^{\prime}=g(\hat{s}^{\prime})$ with $g$ as a decoder from latent states back to observations) is that they may compel the model to reconstruct aspects of the environment that are either inherently unpredictable or irrelevant to task performance. Examples often cited include fine-grained visual details, inconsequential events, or out-of-scene content, which may mislead the model into learning unstable or spurious correlations. Proponents of encoder-only architectures thus suggest that by avoiding this reconstruction step, the resulting WM can focus more selectively on predictable and task-relevant elements. While this motivation is understandable, it remains unclear whether removing the decoder is the effective remedy. In such an architecture as JEPA, supervision occurs solely in the latent space rather than observation space, trading off challenges of pixel-level variability for the risk of indefinability: the predicted latents are not directly grounded in observable data, which makes it difficult to diagnose whether the model is learning meaningful dynamics or collapsing to trivial solutions, an issue we discuss formally in §4.4§4.4$\S$ 4.4§ 4.4.
@@ 4.4 Objective: Learning in Data Space, or Latent Space?
If we construct such a degenerate solution $(h^{*},f^{*})$, this solution satisfies $\mathcal{L}_{\text{latent}}(h^{*},f^{*})=0$, which is a global minimum as $\mathcal{L}_{\text{latent}}(h,f)\geq 0$ for all $(h,f)$. ∎
@@ 5.2 The PAN Architecture
Overall, with its hierarchical, multi-level, and mixed representation architecture, and an encoder-decoder pipeline that threads perception $o$, action $o$, belief $\hat{s}_{i}$, simulated belief $\hat{s}^{\prime}_{i}$, and simulated worlds $o^{\prime}$, PAN represents a general-purpose generative model for simulating actionable real-world possibilities for an agent to perform purposeful reasoning, as we will briefly elude to in §$\S 5.4$. PAN does not sidestep the variability in raw perceptual input, but instead modularizes and organizes it. This enables richer internal simulation of every layer of experience for more powerful agent reasoning and planning.
@@ 5.4 Towards Agentic Reasoning with PAN
Recall in §$\S 2$ we outlined an agent architecture for simulative reasoning using the world model. PAN fits naturally into this paradigm, functioning not merely as a video generator, but as a rich internal sandbox for simulation, experimentation, and foresight.
@@ Appendix A Proof for Theorem 1
Case 1 (Modality Tokenizer): Take $N=T$, and construct a vocabulary $\mathcal{V}$ with size $M_{\epsilon}=\lceil\sqrt{TD}\tilde{K}\tilde{\epsilon}^{-1}\rceil^{D}$. We then define the encoding function for $x$ as $f_{\epsilon}(x)=(c_{1},\dots,c_{T})$ where $c_{t}\in\mathcal{V}$. We define the index of $c_{t}$ as a $D$-digit number of base-$M_{\epsilon}$
@@
We map each $x_{td}$ to a number $u_{td}\in[0,1]$ represented by $L$ digits of base-$N$, where the $n$-th digit $d_{n}(u_{td})=c_{s+n}$. In this manner, $x_{td}$ will correspond to subsequence $(c_{s+1},c_{s+2},\dots,c_{s+L})$, where $s=(td-1)\lceil\log_{M}\sqrt{TD}\tilde{K}\tilde{\epsilon}^{-1}\rceil$ and $L=\lceil\log_{N}\sqrt{TD}\tilde{K}\tilde{\epsilon}^{-1}\rceil$. Thus, we have that
@@ Appendix C Proof for Proposition 2
Hence, the total loss under $(\tilde{h},\tilde{f})$ is strictly less than under $(h’,f’)$ ’ ):
@@ Appendix D Proof of Theorem 2
Latent Prediction $\text{Latent Prediction }p_{h\circ f}:$ $\hat{s}^{\prime}\mid o,a\sim\mathcal{N}(f(h(o),a),I)$
Latent Target $\text{Latent Target }\hat{q}_{h}:$ $\hat{s}^{\prime}\mid o^{\prime}\sim\mathcal{N}(h(o^{\prime}),I)$
Observation Prediction $\text{Observation Prediction }p_{h\circ f\circ g}:$ $\hat{o}^{\prime}\mid o,a\sim\mathcal{N}(g\circ f(h(o),a),I)$
@@
Roundtrip Latent Prediction $\text{Roundtrip Latent Prediction }p_{h\circ f\circ g\circ h}:$ $\hat{s}^{\prime}\mid o,a\sim\mathcal{N}(h\circ g\circ f(h(o),a),I).$

### v3

@@ 4.3 Architecture: Autoregressive Generation is Not the Enemy
A common concern with encoder-decoder architectures (which defines an additional function $\hat{o}^{\prime}=g(\hat{s}^{\prime})$ with $g$ as a decoder from latent states back to observations) is that they may compel the model to reconstruct aspects of the environment that are either inherently unpredictable or irrelevant to task performance. Examples often cited include fine-grained visual details, inconsequential events, or out-of-scene content, which may mislead the model into learning unstable or spurious correlations. Proponents of encoder-only architectures thus suggest that by avoiding this reconstruction step, the resulting WM can focus more selectively on predictable and task-relevant elements. While this motivation is understandable, it remains unclear whether removing the decoder is the effective remedy. In such an architecture as JEPA, supervision occurs solely in the latent space rather than observation space, trading off challenges of pixel-level variability for the risk of indefinability: the predicted latents are not directly grounded in observable data, which makes it difficult to diagnose whether the model is learning meaningful dynamics or collapsing to trivial solutions, an issue we discuss formally in §4.4$\S$ 4.4§ 4.4.
@@ 4.4 Objective: Learning in Data Space, or Latent Space?
[…]
@@ 5.2 The PAN Architecture
Overall, with its hierarchical, multi-level, and mixed representation architecture, and an encoder-decoder pipeline that threads perception $o$, action $o$, belief $\hat{s}_{i}$, simulated belief $\hat{s}^{\prime}_{i}$, and simulated worlds $o^{\prime}$, PAN represents a general-purpose generative model for simulating actionable real-world possibilities for an agent to perform purposeful reasoning, as we will briefly elude to in §5.4$\S$ 5.4§ 5.4. PAN does not sidestep the variability in raw perceptual input, but instead modularizes and organizes it. This enables richer internal simulation of every layer of experience for more powerful agent reasoning and planning.
@@ 5.4 Towards Agentic Reasoning with PAN
Recall in §2$\S$ 2§ 2 we outlined an agent architecture for simulative reasoning using the world model. PAN fits naturally into this paradigm, functioning not merely as a video generator, but as a rich internal sandbox for simulation, experimentation, and foresight.
@@ Appendix A Proof for Theorem 1
[…]
@@
[…]
@@ Appendix C Proof for Proposition 2
[…]
@@ Appendix D Proof of Theorem 2
$\text{Latent Prediction }p_{h\circ f}:$ $\hat{s}^{\prime}\mid o,a\sim\mathcal{N}(f(h(o),a),I)$
$\text{Latent Target }\hat{q}_{h}:$ $\hat{s}^{\prime}\mid o^{\prime}\sim\mathcal{N}(h(o^{\prime}),I)$
$\text{Observation Prediction }p_{h\circ f\circ g}:$ $\hat{o}^{\prime}\mid o,a\sim\mathcal{N}(g\circ f(h(o),a),I)$
@@
$\text{Roundtrip Latent Prediction }p_{h\circ f\circ g\circ h}:$ $\hat{s}^{\prime}\mid o,a\sim\mathcal{N}(h\circ g\circ f(h(o),a),I).$